OpenStudy (anonymous):

3x - y = 12 x + y = 8 What is x?

OpenStudy (anonymous):

3x-y=12 +y +1 3x=13 4.33333333333333333333333333333333

OpenStudy (anonymous):

@DaWizjr haha no way. add both the equations. 3x-y = 12 + x + y = 8 We get 4x = 20 and hence x = 5

OpenStudy (anonymous):

^ LOL

OpenStudy (tkhunny):

Hey, @thabiggestboss47, I haven't seen you do any work, yet. So far, you've only gotten others to solve things for you. Let's see YOU do one. Show all your work.

OpenStudy (anonymous):

thank you luckythebest . and as for DaWizjr you are a complete dumbass. oh yeah tkhhunny how bout you mind your business man.

OpenStudy (anonymous):

flutter you^ not my fault you slow

OpenStudy (tkhunny):

@thabiggestboss47 1) You put problems on a public board hoping to receive assistance. 2) Volunteers decide what kind of assistance you need. 2a) Some volunteers think you need your homework done for you. 2b) Some volunteers want to see you learn something. 3) In any case, the interaction of volunteer with original poster is generally the choice of the volunteer. 4) Having made the public appeal, the original poster doesn't really get to dictate what sort of assistance is provided. To sum up, since my interest is in you actually learning the material with which you have been tasked, my comments are my business - even on a question posted by you. The more you want to learn, the more you show your work. If I am the only one demanding that you show your work, then your learning requires you to listen to me.

OpenStudy (anonymous):

^